ACTIVITY 8. APPLYING HYPOTHESIS TESTING

Solve the problem:

A recent survey says that Filipino children spend an average of 4 hr a day playing online games with a standard deviation of 30 minutes. A random sample of 9 children is taken from a normally distributed population of children who spend an average of 4 hr and 30 min playing online games. Using the 1% level of significance, would you conclude that the statement given in the survey is correct?


1
Expert's answer
2022-05-03T10:59:08-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=4"

"H_a:\mu\not=4"

This corresponds to a two-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha = 0.01," and the critical value for a two-tailed test is"z_c = 2.5758."

The rejection region for this two-tailed test is "R = \{z: |z| > 2.5758\}."

The z-statistic is computed as follows:


"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{4.5-4}{0.5/\sqrt{9}}=3"

Since it is observed that "|z| = 3 >2.5758= z_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value is "p =2P(Z>3)=2(0.00135)=0.0027," and since "p = 0.0027 < 0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"  is different than 4, at the "\alpha = 0.01" significance level.